Fermi National Laboratory
 

Operation Summary for May 20, 2011 - May 23, 2011

 

The Integrated Luminosity for the period from 5/16/11 to 5/23/11 was 59.22 pb-1 (inverse picobarns). NuMI took 4.73E16 of beam during this same time period.

Friday May 20

The day shift began with Tevatron (TEV) personnel conducting hollow beam studies, with Operations monitoring TEV store 8749 (87.42E30) and the Recycler stash (366.0E10), with the Antiproton Source (Pbar) stacking (13.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off due to target replacement, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, TEV personnel concluded the hollow beam studies at 11 AM. Operations terminated store 8749 and prepared CDF and D0 for collision hall accesses.

During the evening shift, Operations put the TEV into HEP shot setup at 4:25 PM and established store 8753 at 5:25 PM with an average initial luminosity (AIL) of 397.1E30.

Saturday May 21

The midnight shift began with Operations monitoring the store 8753 (148.39E30) and the stash (313.7E10), with Pbar stacking (27.7E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8753 at 7:25 AM and put the TEV into shot setup.

During the day shift, Operations established store 8754 at 8:40 AM with an AIL of 376.0E30.

During the evening shift, Operations terminated store 8754 at 7:40 PM. Operators reported that just before store termination, the TEV sector B1 LCW pump had tripped off and a TEV LCW leak alarm posted. Operators searched the TEV service building for the source of the LCW leak, but found nothing. Operators reported that the B1 LCW pump would not reset. The FESS Duty Electrician reported that he could not close B1 LCW pump breaker. A TEV LCW person told operators to turn on the B3 LCW and pond pumps. TEV water personnel later came in and made up water to the Central Utility Building (CUB) spheres.

Sunday May 22

The midnight shift began with operators checking the TEV orbits, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations put the TEV into shot setup at 12:05 AM, but the TeV quenched during the second transfer of antiprotons. Operators investigated and found that a D0 quadrupole power supply (D0QT6) regulator’s output had wandered. Operators reported that during the change out of the regulator, one of the load cables fell out of its connector. Operators called in an EE Support technician and he replaced the connector. The TEV Coordinator came in and determined that the quench occurred milliseconds before the antiproton beam was injected. Only one polarity of the E48 kicker fired when the antiprotons entered the TeV. Operators recovered the TEV and started a new shot setup at 7:10 AM.

During the day shift, Operations established store 8759 at 8:06 AM with an AIL of 363.9E30. Operators reported that a Booster RF station (BFR13) tripped off and would not reset; operators turned on BRF19 in its place. Operations terminated store 8759 at 6:20 AM, put the TEV into shot setup, and established store 8760 at 7:40 PM with an AIL of 386.0E30.

During the evening shift, at 9 PM, a lightning strike caused Recycler kickers KPS2A and KPS3A to pre-fire, which aborted the stash, causing the loss of 78E10 antiprotons. Operators reported that a Recycler cooling amplifier tripped off in the MI tunnel. Operators prepared MI for an access. Operators called in a Recycler cooling technician and he replaced the amplifier. Operators checked the Fluorinert levels on KPS3A. Operators reported that the lightning strike also caused a communication loss with the Electron Cooling system. Operators called in Recycler and Controls personnel; they determined that the fault was between the CAMAC controls crate and the network card connected to the local intranet. Operators reported that due to the loss of communications, the Pelletron was down for the rest of the shift.

Monday May 23

The midnight shift began with Operations monitoring store 8760, with stacking off, and with no beam to MiniBooNE.

During the midnight shift, operators reported that Pbar had resumed stacking and Operations had resumed sending beam to MiniBooNE by 12:02 AM. Operations reported that the Tevatron Electron Lens had tripped off several times during the shift. Operators reported that the F0 pond water supply pressure dropped and the TeV RF watchdog alarmed. Operators reported that MIRF4 tripped off numerous times and eventually they bypassed the station.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow TEV personnel to conduct end-of-store studies and allow Booster personnel to also conduct a study.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/23/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory